OpenStudy (anonymous):

Could somebody help me to derivate this g(w)= sin (cos 3w)- 3 cos^2 2w

OpenStudy (anonymous):

cos ( cos (3w))*(-sin(3w))*3-

OpenStudy (anonymous):

6cos(2w)*(-sin(2w))*2

OpenStudy (anonymous):

-3 cos (cos (3w)) sin(3w)+12cos(2w)sin(2w)

OpenStudy (anonymous):

* sin (2x)=2 sin x cos -3 cos (cos (3w)) sin(3w)+6 sin(4w)

OpenStudy (irishboy123):

do each term separately and then add the up at the end. taking them in turn: : g = sin (cos 3w) you might wish to break this up a bit more to make it easier. you could say that g = sin (x) where x = cos (3w) then use chain rule: dg/dw = dg/dx * dx/dw : g = - 3 cos^2 2w this is easier but again, if you find it daunting, you might try doing the same kind of thing. you can say g = - 3 x^2, where x = cos (2w) chain rule it: dg/dw = dg/dx * dx/dw

OpenStudy (anonymous):

@IrishBoy123 i don't get very well could explain me how to solve g=-3cos ^2 2w

OpenStudy (irishboy123):

if g=-3cos ^2 2w then let x = cos 2w thus g = -3 x^2 yes?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

what are you looking for the derivative of?

OpenStudy (misty1212):

it is this? \[g(w)= \sin (cos (3w))- 3 \cos^2 (2w)\]

OpenStudy (anonymous):

yeah

OpenStudy (misty1212):

gotta use the chain rule all the way down here

OpenStudy (misty1212):

first lets look at \[\sin(\cos(3w))\] the derivative of sine is cosine so start with \[\cos(\cos(3w)\times \] then the derivative of cosine is minus sine so \[\cos(\cos(3w)\times (-\sin(3w))\times \]

OpenStudy (misty1212):

and finally the derivative of \(3w\) is \(3\) final answer \[\cos(\cos(3w)\times (-\sin(3w))\times3\]

OpenStudy (misty1212):

clean it up a little and write \[-3\cos(\cos(3w))\sin(3w)\] and is just the first part

OpenStudy (misty1212):

you good with that? we still have to do \[-3\cos^2(2w)\]

OpenStudy (anonymous):

In that part is where I got confuse because is square

OpenStudy (misty1212):

\[\frac{d}{dx}f^2(x)=2f(x)f'(x)\]

OpenStudy (misty1212):

the derivative of something squared is two time something, times the derivative of something

OpenStudy (misty1212):

in your case the something is \(\cos(2w)\) and the derivative of that is \[2\sin(2w)\]

OpenStudy (misty1212):

the derivative of the second part is therefore \[+12\times \cos(2w)\sin(2w)\]

OpenStudy (anonymous):

then whats is next S:?